### v4

# Belief Bias Identification

Abstract
This paper proposes a unified theoretical model to identify and test a comprehensive set of probabilistic updating biases within a single framework. The model achieves separate identification by focusing on the updating of belief distributions, rather than point beliefs alone. Estimating the model in a laboratory experiment reveals significant individual heterogeneity: all tested biases are present and exhibit systematic co-occurrence patterns across individuals, with motivated-belief biases (optimism and pessimism) and sequence-related biases (gambler’s and hot-hand fallacy) emerging as key drivers of biased inference. At the population level most biases average out, but base-rate neglect remains a persistent influence. This study contributes to the belief-updating literature by providing a methodological toolkit for researchers examining links between conflicting biases and connections between updating biases and other behavioral phenomena.
00footnotetext: For their invaluable support, I would like to thank Elias Tsakas, Thomas Meissner and Matthias Wibral. For helpful comments, I thank three anonymous referees, Mohammed Abdellaoui, Miguel Ballester, En Hua Hu, Peiran Jiao, Emmanuel Kemel, Diego Marino-Fages, Antonio Penta, Arno Riedl, David Rojo Arjona, Michael Thaler, Severine Toussaert, and seminar participants at IMEBESS in Riga, ASFEE in Grenoble, Behavioural Game Theory Workshop in Norwich, Cognitive Foundations of Decision Making in Rabat, PhD conference in Micro Theory and Behavioral Economics in Berlin, and Maastricht University. I acknowledge funding from the Maastricht Graduate School of Business and Economics (GSBE) for primary data collection No. G.23.5168. Homepage: pgonzalezfernandez.com. E-mail: pedro.gonzalez-fernandez@awi.uni-heidelberg.de.
Keywords: Belief Updating, Belief Biases, Probabilistic Reasoning
JEL codes: D01, D90
1 Introduction
Over the last decades, substantial effort has gone into identifying and categorizing belief-updating biases, both within Economics and Psychology (Edwards, 1968; Tversky and Kahneman, 1974; Grether, 1980). Scholars have typically studied these biases separately. As a result, it has become difficult to tell whether a person’s updating behavior should be attributed to one bias or another. Despite recent progress toward more comprehensive representations of biased updating (Benjamin, 2019; Stango and Zinman, 2023; Bordalo et al., 2023, 2025), there is still a lack of a unifying framework that can separate multiple potentially conflicting biases within the same empirical setting.
The basic problem is that different belief-updating biases may be observationally equivalent. Consider a person who appears to update “too much” after receiving new information. A traditional model may ask whether this reflects base-rate neglect, namely underweighting prior beliefs, or overinference, meaning placing too much weight on the new signal. But other biases may generate a similar updating pattern. If the person has preferences over outcomes, is she overreacting because she generally overinterprets information, or because she is optimistic or pessimistic about the state she prefers? Is she instead reacting strongly because the signal confirms her prior beliefs? Or does the same posterior reflect unusually concentrated beliefs, that is, overprecision? Once some biases are left outside the model, behavior may be attributed to the wrong one. In that sense, some biases may appear empirically important partly because other competing distortions are not being modeled.
Distinguishing such confounded biases matters for at least two reasons. First, different biases can lead to very different actions even when they produce similar posteriors. For example, a person exhibiting confirmation bias may keep voting for the same political party after receiving social-media information that aligns with her previous beliefs, while someone who simply neglects prior information might be more likely to change her mind if reminded of the actual base rate. Second, once these biases are separately identified, one can ask which distortions are more prevalent, which matter most for inference, and which ones tend to appear together within individuals.
A central challenge in addressing this problem is methodological. The way beliefs are usually measured may not be rich enough to identify a broad set of potentially confounded biases. Much of the literature relies on point beliefs to study biased updating (Benjamin, 2019). While convenient, point beliefs offer limited flexibility to distinguish between different updating patterns. This paper proposes an alternative framework based on belief distributions.
The paper constructs a structural model that allows multiple conflicting belief-updating biases to be separately identified using prior and posterior belief distributions. [^1] While Jiao et al. (2020) use belief distributions to incorporate several biases simultaneously, their framework is not well suited to probabilistic biases, relies on normally distributed beliefs, incorporates a more limited set of distortions, and is not taken to experimental data. Belief distributions can be understood either as beliefs over multiple outcomes or as uncertainty over one’s own beliefs when the underlying state space is binary, in a sense related to cognitive uncertainty (Enke and Graeber, 2023). In either interpretation, the key idea is that belief distributions provide a sharper language for defining and distinguishing belief biases.
The theoretical framework considers an agent who observes a sequence of Bernoulli trials. Prior beliefs are assumed to follow a beta distribution and, under Bayesian updating, posteriors remain beta distributed because the beta distribution is conjugate to the binomial distribution. [^2] This updating environment is comparable to settings used in the ambiguity-aversion literature (see Abdellaoui et al. (2025) for an example) and in work on cognitive uncertainty (Enke and Graeber, 2023). To introduce biases, the model allows for distortions of the likelihood and the prior. These distortions generate non-Bayesian posterior beta distributions and make it possible to identify core biases such as over- and underinference (Khaw et al., 2021; Ba et al., 2022; Augenblick et al., 2025b) and base-rate neglect or overuse (Benjamin et al., 2019; Bucher and Glimcher, 2022; Enke et al., 2023). The framework also accommodates asymmetries in reactions to good and bad news, preference-based biases (Eil and Rao, 2011; Zimmermann, 2020; Möbius et al., 2022), confirmation bias (Rabin and Schrag, 1999; Charness and Dave, 2017; Zhenxun, 2024), sequence-related biases such as the hot-hand and gambler’s fallacies (Rabin, 2002; Rabin and Vayanos, 2010; Prat-Carrabin et al., 2024), and precision distortions (Moore et al., 2015; Augenblick et al., 2025a) that affect the variance of posterior beliefs.
The model is estimated using data from a novel laboratory experiment. Participants solve a series of urn tasks in which they must guess the percentage of red balls in a selected urn drawn from a pool of 99 urns with varying red-blue compositions. They receive information through sequences of ball draws from the selected urn and report their belief beta distributions twice: once after an initial signal, and again after observing a second sequence of independently drawn signals from the same urn.
The results support the idea that richer models of biased updating reduce the apparent importance of some biases found in more parsimonious specifications. At the population level, overinference and base-rate neglect are both significant when the model allows only these distortions. Once the model incorporates a broader set of biases, however, the evidence for overinference disappears and base-rate neglect remains as the only distortion at the aggregate level. This aggregate bias distribution is also broadly consistent with the classic reduced-form belief-updating literature when the same environment is translated into a traditional binary-state specification.
At the same time, aggregate results mask substantial individual heterogeneity. [^3] The importance of individual heterogeneity for biased updating has recently been emphasized by Khaw et al. (2021) and Alós-Ferrer and Garagnani (2023). Their work supports the idea that average reports may look almost Bayesian or noisy, while individual-level data reveal systematic but heterogeneous deviations that partly cancel out in the aggregate. The individual-level analysis is therefore more informative. First, all tested biases are present to some extent in the data once the full array of biases is taken into account, revealing a much richer pattern of distorted updating than is visible in the aggregate. This also makes it possible to identify distorted behavior that would otherwise be statistically indistinguishable from Bayesian updating or too noisy to classify. Second, motivated-belief biases (optimism and pessimism) and sequence-related biases (the hot-hand and gambler’s fallacies) emerge as the main drivers of biased inference, with the hot-hand fallacy standing out as the most commonly exhibited distortion, while confirmatory biases are comparatively scarce. Finally, biases do not appear independently. Instead, they exhibit systematic co-occurrence patterns across individuals. Once again, sequence-related biases are central: A group of distortions (especially overprecision, overinference, and motivated-beliefs in general) seem to cluster around the hot-hand fallacy, while a different class of biases (such as prior-based distortions) tend to be jointly exhibited with the gambler’s fallacy.
The paper contributes to the literature in three ways. First, it contributes to recent work that seeks to understand the links between different updating distortions. For example, Heger and Papageorge (2018) and Gneezy et al. (2023) study how wishful thinking can affect overconfidence; Charness and Dave (2017) and Zhenxun (2024, 2023) distinguish behavior stemming from motivated beliefs and unmotivated confirmation bias; Aydogan et al. (2025) build on Rabin and Schrag (1999) to separate conservatism from confirmation bias; and Chopra et al. (2024) study the trade-off between accuracy concerns and belief-confirmation motives in political information processing. These are important steps toward understanding how biases interact. More generally, however, in the absence of a framework that accommodates a broad range of distortions simultaneously, empirical links between biases may themselves be confounded. Bias $A$ may appear related to bias $B$ only because a third bias $C$ is omitted.
Second, the paper contributes to research that seeks belief-based explanations for behavioral phenomena in settings where several competing distortions are plausible. Political polarization, for instance, has been explained from the perspective of overconfidence (Ortoleva and Snowberg, 2015), which is closely related to overprecision, and also from the perspective of confirmation bias (Del Vicario et al., 2017). In finance, the disposition effect has been linked to motivated beliefs (Heinke et al., 2023), the gambler’s fallacy (Jiao, 2017), and general underinference (Pitkäjärvi, 2022). Other examples include work linking confirmation bias to stylized facts in financial markets (Pouget et al., 2017), or confidence distortions to poor investment performance (Ahmad and Shah, 2020) and biased memory (Huffman et al., 2022). A framework that distinguishes conflicting biases more cleanly can therefore help identify which distortions are actually driving particular behavioral regularities, and whether some candidate explanations are partly artifacts of less complete models.
Third, the paper contributes to the literature on behavioral interventions. Boosting refers to de-biasing techniques that aim to improve decision-making when individuals suffer from cognitive distortions. One challenge emphasized in this literature is that interventions often target one bias at a time, even though several may be jointly at work (Kahneman et al., 2021). A method that distinguishes between multiple conflicting belief-updating biases can therefore help direct interventions toward the distortions that are most prevalent or that matter most for inference.
The remainder of the paper is organized as follows. Section 2 introduces the theoretical framework and progressively incorporates different biases into the model. Section 3 presents the experimental design and the belief-measurement tool. Section 4 brings the theoretical and experimental components together and compares two regression models that differ in the number of biases they incorporate. Section 5 discusses the results, and Section 6 concludes.
2 A model of multiple belief biases
2.1 Theoretical framework
Let an agent observe a signal $S=\{s_{1},...,s_{n}\}$ consisting of the realization of $n$ independent and identically distributed Bernoulli trials. Let $p$ denote the probability of success of each trial ($s_{i}=1$) and 1p1-p denote the probability of each failure ($s_{i}=0$). For such data generating process (DGP), the likelihood function is the probability mass function of a binomial distribution with parameters $(n,p)$:
$L(p|s_{1},\dots,s_{n})=\binom{n}{k}p^{k}(1-p)^{n-k}$ (1)
where $k={\sum_{i=1}^{n}s_{i}}$ and (nk)(n-k) are the number of successes and failures in the DGP respectively.
Let $\Omega=(0,1)$ be the set of possible values that $p$ may take, [^4] The state-space set $\Omega$ can have two different interpretations. One could either consider an agent who is not uncertain about her own beliefs, or one who is indeed uncertain about her own beliefs (à la cognitive uncertainty (Enke and Graeber, 2023)). In the former case, the set $\Omega$ would specify the state-space of an agent who forms beliefs over every possible realization of the objective parameter $p\in(0,1)$. In the latter case, the agent would form beliefs over a binary state-space (whether the signal $s_{i}$ takes value $1$ or 0), and the set $\Omega$ would, in this case, represent the subjective state-space ($p$), which describes the set of possible states where the agent expresses uncertainty about her own beliefs. The model is agnostic to either one of these interpretations. and let a prior belief $\pi(p)$ be beta distributed with prior parameters $(a_{0},b_{0})$. Namely,
$\pi(p|a_{0},b_{0})=\frac{1}{B(a_{0},b_{0})}p^{a_{0}-1}(1-p)^{b_{0}-1}$ (2)
where $a_{0}$, $b_{0}$ >0>0, and $B(.)$ is the beta function.
Given a prior $\pi(p)$ and a likelihood $L(p|s_{1}...s_{n})$ the agent forms a posterior. As the beta distribution is a conjugate prior of the binomial distribution, a Bayesian agent updates her beliefs such that her posterior distribution of $p$, $\pi(p|a_{n},b_{n})$, is also beta distributed with parameters $a_{n},b_{n}$. This means that: [^5] See Appendix A for a proof; as well as a proof to derive eq.(4) and (5), which suffice to derive all other equations of posterior parameters in section 2.
$a_{n}=k+a_{0}\qquad b_{n}=n-k+b_{0}$ (3)
In order to incorporate updating biases, let a non-Bayesian agent use a distorted likelihood and prior when she updates. [^6] These distortions are often seen (Benjamin, 2019) as part of an “as-if model.” This means that the model does not take the stand that biased agents actually follow Bayes’ Theorem with different likelihood and prior functions, but instead, that these distortions imply equivalent behavior to agents interpreting too little, or too much, from information signals (or prior beliefs) when they update. These distortions can be expressed as exponential deviations of the likelihood and prior. Namely, $\tilde{L}(p|s_{1},\dots,s_{n})=(L(p|s_{1},\dots,s_{n}))^{\gamma}$ would represent the distorted likelihood and $\tilde{\pi}(p)=(\pi(p))^{\delta}$ would represent a distorted prior of the non-Bayesian agent ($\gamma,\delta>0$). The parameters $\gamma\neq 1$ and $\delta\neq 1$ indicate deviations from Bayesian updating due to distortions of the likelihood and prior respectively. With these modified functions such a non-Bayesian agent has a posterior beta distribution with parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
$\tilde{a}_{n}=\gamma k+\delta(a_{0}-1)+1$ (4)
$\tilde{b}_{n}=\gamma(n-k)+\delta(b_{0}-1)+1$ (5)
In particular $\gamma>1$ indicates “overinference” while $\gamma\in[0,1)$ shows “underinference”; that is, believing that the information signal $S$ is more/less informative than it actually is. A parameter of $\gamma<0$ reflects updating against the information signal. Similarly, parameter $\delta>1$ indicates “base-rate overuse” and $\delta<1$, “base-rate neglect”; which implies that the prior is more/less informative than a Bayesian agent perceives it to be. This bias structure conceptually resembles Grether-regressions, [^7] That is, binary-state regressions where inference biases and base-rate biases are identified. as the same type of biases can be identified. The next section shows how equations (4) and (5) can be extended to incorporate incrementally more biases.
2.2 Introducing multiple belief biases
2.2.1 Asymmetries between successes and failures & Preference-based biases
Equations (4) and (5) assume that deviations from the Bayesian agent in the likelihood are symmetric for successes and failures. That is, the agent over or under-reacts to “positive information” the same way she over or under-reacts to “negative” information. For a non-Bayesian agent this need not be the case. Consider instead that the agent weights successes and failures of the data generating process differently. Then her likelihood function would be $\tilde{L}(p|s_{1},\dots,s_{n})\propto p^{\alpha k}(1-p)^{\beta(n-k)}$. In turn, her posterior would be beta distributed with shape parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
$\tilde{a}_{n}=\alpha k+\delta(a_{0}-1)+1$ (6)
$\tilde{b}_{n}=\beta(n-k)+\delta(b_{0}-1)+1$ (7)
where $\alpha\neq\beta$ indicates asymmetric reactions to successes and failures in the DGP. Equations (6) and (7) are especially interesting if the agent has preferences over the state-space. Suppose this is the case, and suppose further that preferences are expressed by a utility function, which is continuous and monotonically increasing over $p$. Then, every success $k$ would be informative of a higher realization of $p\in(0,1)$, i.e. a preferred state. Conversely, every failure (nk)(n-k) is informative of a lower realization of $p$. Therefore, successes and failures can be interpreted as pieces of good and bad news respectively. This means that $\alpha>1$ (overreacting to positive information) or $\beta<1$ (underreacting, or updating against the signal when confronted with negative information) can be interpreted as optimism bias; while $\alpha<1$ or $\beta>1$ would indicate pessimism. Furthermore, the model also captures motivated asymmetries: $\alpha>\beta$ represents the good news effect, while $\alpha<\beta$ implies there is bad news effect. [^8] Asymmetries between successes and failures can also be present at the level of the prior. These distinctions are introduced in section 4, but left out of the main theoretical discussion. This is without loss of generality as the resulting equations are identical to equations (9) and (10) except for having separate $\rho$ and $\delta$ parameters for successes and failures. (see section 4.2)
2.2.2 Confirmation Bias
Confirmation bias is modeled as a positive correlation between overreaction to information signal $S$, and how confirming the signal is. The degree of confirmation of the signal is unrelated to the agent holding preferences [^9] A discussion about the distinction between motivated and unmotivated confirmation bias can be found in Zhenxun (2023) over the state-space. In particular, the degree of confirmation $c$ is expressed as the area, in the density function of the prior, comprised between the expected value of the prior $E(\pi(p))$ and the mean of the information signal $k/n$. Formally:
$c=\bigg|\int_{(k\pm\varepsilon)/n}^{E(\pi(p))}\pi(p|a_{0},b_{0})\,dp\bigg|$ (8)
Equation (8) specifies a relative measure of confirmation [^10] In equation (8), the signal mean $k/n$ is replaced by $(k\pm\varepsilon)/n$. This is because $p$ is not strictly defined at $p=0$ and $p=1$. Thus, one must take $(k+\varepsilon)/n$ as the inferior limit if $k=0$; and $(k-\varepsilon)/n$ if $k=n$. (see Figure 1 for an example). The higher the value of $c$, the less confirming a signal will be. To incorporate over or under reaction to confirmation a separate term $\rho c$ is added to the distorted likelihood function. Equations (6) and (7) are modified as follows:
$\tilde{a}_{n}=\alpha k+\rho c+\delta(a_{0}-1)+1$ (9)
$\tilde{b}_{n}=\beta(n-k)+\rho c+\delta(b_{0}-1)+1$ (10)
where $\rho<0$ indicates confirmation bias (as $c$ becomes smaller the perception of the number of successes or failures grows); and $\rho>0$ indicates disconfirmation bias.
Figure 1: Example of confirmation measure for an updating problem with $a_{0}=b_{0}=9$, prior mean $E(\pi(p))=0.5$ and information signal of 8 successes and 2 failures ($k/n=8/10$)
2.2.3 Sequence-related biases and unrelated inference biases
Overreaction or underreaction to information signals can also be driven by other reasons different to the presence of preferences over the state-space. I distinguish between: sequence-related biases (e.g., the hot-hand fallacy and the gambler’s fallacy) and baseline inference. That is overreaction to information signals when preferences over the state-space are not at stake.
Sequence-related biases: Consider a partition of the signal space $S$. Namely, $S_{1}=\{s_{1}\dots s_{m}\}$ and $S_{2}=\{s_{m+1}\dots s_{n}\}$ where $n>m$. Suppose all of the $s_{i}\in S_{2}$ are either successes (i.e. $\sum_{n-m}^{n}s_{i}=n-m$, $s_{i}\in S$) or failures (i.e. $\sum_{n-m}^{n}s_{i}=0$, $s_{i}\in S$). In this context, the hot-hand fallacy is defined as inferring too much from information signals after observing the last nmn-m consecutive successes (i.e. $\alpha>1$) or failures (i.e. $\beta>1$). Conversely, gambler’s fallacy is defined as inference against the information signal after observing a the last nmn-m consecutive successes (i.e. $\alpha<0$) or failures (i.e. $\beta<0$). Thus, sequence-related biases are modeled as a context-dependent component of inference distortions, rather than as a separate channel.
Inference without preferences (baseline inference): Consider an almost equivalent state-space set, $\Omega_{NP}=(0,1)$ over which the agent form beliefs but where, contrary to $\Omega$, the agent does not hold any preference. So any element $\hat{p}$ of the set $\Omega_{NP}$, will have the characteristic that $u(\hat{p})=0$. Because this is the only difference between sets, only $\alpha$ and $\beta$ coefficients in equations (9) and (10) differ between forming beliefs over $\Omega$ or $\Omega_{NP}$. Therefore, when the agent forms beliefs over $\Omega_{NP}$, $\alpha$ and $\beta$ coefficients can be interpreted as over(under) inference for successes and failures, independently of preferences over the state-space.
The biases in this subsection are only separably identified when the agent faces multiple belief-elicitation decisions that generate variation both in whether preferences are at stake (tasks in $\Omega$ versus $\Omega_{NP}$) and in whether the observed signal contains a streak.
2.2.4 Under and overprecision
Precision biases (overprecision and underprecision) are those which are strictly related to the variance of the agent’s posterior distribution in relation to the Bayesian variance. [^11] As an alternative to capture deviations in the second moment of the distribution, one could also use the concentration parameter of the Bayesian posterior ($\kappa_{n}=a_{n}+b_{n}$) to define under and overprecision. This alternative definition has the advantage of isolating precision biases within equations (9) and (10). (i.e. Underprecision: $\alpha+\beta<2$, and Overprecision: $\alpha+\beta>2$). However, it imposes harsh restrictions on $\alpha$ and $\beta$ parameters, making precision biases harder to be detected. Underprecision implies that the agent’s overall distribution of posterior beliefs is more dispersed than that of a Bayesian agent. Conversely, overprecision yields a posterior distribution which is less dispersed over the values of $p$ than the distribution a Bayesian agent would have. That is:
$\tilde{Var_{n}}=\nu\times Var_{n}$ (11)
where $\tilde{Var_{n}}$ is the variance of the agent’s posterior beta distribution, and $Var_{n}$ is the Bayesian variance. In equation (11), $\nu<1$ indicates overprecision, while $\nu>1$ indicates underprecision.
Summary.
For ease of reference, Table 1 in Section 4.2 summarizes the mapping from parameter restrictions to the bias labels used throughout the paper.
3 Experimental Design
The experiment provides a setting in which the belief biases described in the theoretical framework of Section 2 can be identified at both the individual and population levels. It was conducted at the behavioral and Experimental Economics Laboratory (BEELab) at Maastricht University. A total of 88 participants were recruited, and each completed 30 belief-elicitation tasks. In each task, participants reported their belief distributions twice, after observing two consecutive information signals. The average payment per participant was 15.9 euros. The experiment was pre-registered in October 2023.121212See the preregistration at https://aspredicted.org/285v-b8cg.pdf The full instructions of the experiment can be found in Appendix F.
3.1 The belief-elicitation task
In each of the 30 tasks, participants observed a pool of 99 urns, each containing 100 balls. Each urn contained a different composition of red and blue balls. Thus, Urn 1 contained 1 red ball and 99 blue balls, Urn 2 contained 2 red balls and 98 blue balls, and so on, up to Urn 99, which contained 99 red balls and 1 blue ball. One of these urns was then selected at random from the pool, but its content was not revealed to participants. Their task was to guess the percentage of red balls in the selected urn (see Figure 2). To do so, participants received information signals by observing two sequences of balls drawn with replacement. In the first sequence, either one, two, or three balls were drawn at random from the selected urn.
Figure 2: Urn selection
After observing this first sequence, participants were asked to report their belief distribution [^13] See Section 3.2 for details about the elicitation of beta distributions. (Figure 3(a)) about the percentage of red balls in the selected urn. This first report is treated as a prior. Once this prior distribution was elicited, a second sequence of draws from the same selected urn was shown to participants. In this case, either three, five, or seven balls were drawn at random from the urn. After observing this second information signal, participants were asked to report their posterior belief distribution [^14] Both in the prior and posterior reports, participants needed to update from a given default beta distribution. Before eliciting their prior, the default beta distribution shown was uniform (as this was the exogenously implemented “prior of the prior”). Accordingly, the default beta distribution shown before eliciting the posterior was the participant’s own prior report. (Figure 3(b)). Once this was done, a new urn from the pool of 99 urns was selected, with replacement, and the process was repeated 30 times. All subjects faced the same set of urns and sequences of draws, but the order of tasks was randomized across participants. All belief-elicitation tasks were incentivized using a binarized scoring rule (more details are provided in Section 3.2 and Appendix B).
To assess the role of motivated beliefs, fifteen of the thirty tasks, placed at random positions in the experiment, required participants to report beliefs about urns to which an additional payment was attached. I henceforth refer to these as dollar urns. In these tasks, participants received a payment in cents equal to the (unknown) number of red balls in the selected urn. Thus, participants had a monetary incentive to prefer urns with a higher proportion of red balls.
3.2 Eliciting beta distributions
To elicit belief distributions, I use the tool introduced in Gonzalez-Fernandez et al. (2025), which was validated there against non-parametric elicitation methods. This method is especially convenient in settings where beliefs are expected to follow a parametric form, while remaining closely connected to the broader distribution elicitation literature (Manski, 2004; Goldstein et al., 2008; Harrison et al., 2017; Crosetto and De Haan, 2023).
To report their beliefs, participants were presented with a dynamic graphical interface that allowed them to select their preferred beta distribution (see Figures 3(a) and 3(b)). By moving two sliders, each associated with a different question, participants were able to select a specific beta distribution. These questions were:
1. What percentage of red balls do you expect the selected urn to have?
2. What is your uncertainty level about this percentage?
The first slider, associated with question (1), allows subjects to manipulate the expected value of a beta distribution, while the second slider, associated with question (2), allows them to manipulate its standard deviation. [^15] The probability density function of any beta distribution can also be parameterized by its expected value and variance. Participants were shown a five-minute explanatory video on how to interpret the graph they selected and how to manipulate it. Importantly, while participants adjusted sliders to report their beliefs, the relevant object for the analysis is the belief distribution they constructed, which was later mapped into beta-distribution parameters, rather than the individual slider positions themselves. Subjects were specifically instructed to solve the task graphically, and were also shown that the scale of the graph updates dynamically in order to keep the plot informative. Participants additionally had the option to fix the graph on a constant scale if they wished to do so. Before beginning the relevant belief-elicitation tasks, participants answered related comprehension questions [^16] In order to participate in the experiment, subjects had to answer at least 3 out of 5 comprehension questions correctly. If any comprehension question was answered incorrectly, participants were given a second chance to answer all questions correctly. and were allowed to test the software. For ease of interpretation, the interface did not allow subjects to report bimodal beta distributions. Bayesian reporting is nevertheless always feasible. [^17] Standard ex-post parametric fitting of the beta distribution imposes the same restriction (Engelberg et al., 2009).
More generally, because of the continuous nature of the variance slider, measurement error in uncertainty reports may affect bias identification. Appendix D.1 therefore applies a noise-injection procedure to assess the robustness of the results to measurement error in variance reports. This procedure imposes no restriction on the admissible variance range to rule out bimodality.
(a) Figure 3a: Example of prior belief elicitation after observing a first sequence of draws from a random urn showing “red, red”. This example shows a specific selection of a scaled beta distribution with an expected value of 56 and a standard deviation of 18.47.
(b) Figure 3b: Example of posterior belief elicitation after observing a second sequence of draws showing “blue, red, blue, red, blue”. The updated beta distribution selected in this example has an expected value of 51 and a standard deviation of 7.44.
To incentivize truthful reporting, a binarized scoring rule was implemented in every task. Specifically, I follow a method similar to that suggested by Schlag and Van der Weele (2013), who propose scoring rules to incentivize different moments of a probability distribution (see Appendix B for the specific rule used here). At the same time, I follow Danz et al. (2022) in not explicitly disclosing the exact scoring rule to participants. Instead, participants were told that, in order to maximize their expected payoff, they should always truthfully report both their guess for the percentage of red balls in the selected urn and their uncertainty about that percentage (that is, the mean and standard deviation of the beta distribution they selected).
4 Baseline and Complete model regressions
This section presents the empirical specifications derived from the theoretical framework in Section 2. Section 4.1 introduces a baseline regression that captures only a limited set of belief biases. Section 4.2 then presents the complete specification (equations (14), (15) and (16)), which incorporates the full set of biases studied in the model. Estimating these two specifications side by side makes it possible to assess whether the bias patterns detected in the baseline model remain once richer bias channels are taken into account. Both models are estimated at the population and individual levels, since aggregate estimates may conceal substantial heterogeneity across subjects. Section 4.3 finally introduces a procedure to compare the relative importance of the biases at the individual level. Additionally, Appendix E reports a Grether-style reduced-form specification of the data in the traditional binary-state framework used in the belief-updating literature.
4.1 Baseline model regressions
The baseline model, closely resembles equations (4) and (5). It tests for under(over)-inference and base-rate neglect(overuse):
$\tilde{a}_{n}-1=\gamma_{s}k+\delta_{s}(a_{0}-1)+\varepsilon_{a}$ (12)
$\tilde{b}_{n}-1=\gamma_{f}(n-k)+\delta_{f}(b_{0}-1)+\varepsilon_{b}$ (13)
In equations (12) and (13), successes and failures (variables $k$ and (nk)(n-k)), represent the number of red and blue balls observed in the second sequence of draws provided to participants. Variables $a_{0}$ and $b_{0}$ are the parameters of the beta distribution elicited by subjects after the first sequence of draws, while $\tilde{a}_{n}$ and $\tilde{b}_{n}$ are the parameters of the beta distribution elicited by subjects after observing the second sequence of draws. $\varepsilon_{a}$ and $\varepsilon_{b}$ are the error terms.
The parameter interpretation of equations (12) and (13) is akin to equations (9) and (10). Parameters $\gamma$ and $\delta$ similarly indicate under/over inference and base-rate neglect/overuse respectively. However, equations (12) and (13) differ by acknowledging that these biases ($\gamma_{s}$$\gamma_{f}$) and ($\delta_{s}$$\delta_{f}$) may vary between successes and failures (i.e. realizations of red and blue balls). The presence of those biases is tested by comparing whether the estimated $\gamma$ and $\delta$ parameters are significantly different from their Bayesian values (i.e. $\gamma_{s}=\gamma_{f}=1$ and $\delta_{s}=\delta_{f}=1$).
Note that equations (12) and (13) do not have an intercept. This has a very straightforward theoretical justification. If all the independent variables in those equations take value 0 (i.e. $k=0,n=0,a_{0}=1,b_{0}=1$), both $\tilde{a}_{n}$ and $\tilde{b}_{n}$ take a value of $1$. This means that when one starts updating from a uniform prior distribution ($a_{0}=1,b_{0}=1$), and observes no information signals whatsoever ($k=0,n=0$), one must not update, i.e. remain at such uniform prior after updating ($\tilde{a}_{n}=1,\tilde{b}_{n}=1$). Having an intercept different than zero would imply that agents update in the absence of any kind of information.
4.2 Complete model regressions
In order to test for the presence of all the biases described in section 2.2, I run a slightly modified version of equations (9), (10) and (11). Namely:
$\tilde{a}_{n}-1=(\alpha_{0}+\alpha_{Pref}I_{Pref}+\alpha_{Seq}I_{Seq_{s}})k+\rho_{s}c+\delta_{s}(a_{0}-1)+\varepsilon_{a}$ (14)
$\tilde{b}_{n}-1=(\beta_{0}+\beta_{Pref}I_{Pref}+\beta_{Seq}I_{Seq_{f}})(n-k)+\rho_{f}c+\delta_{f}(b_{0}-1)+\varepsilon_{b}$ (15)
$\tilde{Var_{n}}=\eta+\nu\times Var_{n}+\varepsilon_{v}$ (16)
Variables in equations (14) and (15) are identical to the baseline model with the only exception of variable $c$, the relative measure of confirmation, as described in section 2.2.2. Equation (16) is almost identical to equation (11) in the interpretation of its variables and the $\nu$ parameter indicating over or underprecision ($\nu<1$ vs $\nu>1$). The inclusion of an intercept $\eta$ and an error term $\varepsilon_{v}$ are the only differences. [^18] Similarly to the baseline model, equations (14) and (15) do not have an intercept for the very same reason as outlined in section 4.1.
Parameters $\rho$ and $\delta$ in equations (14) and (15) capture confirmation biases, and base-rate distortions respectively. As in the baseline case, equations (14) and (15) differ by acknowledging that these parameters may differ between successes and failures ($\rho_{s}$ vs $\rho_{f}$) and ($\delta_{s}$ vs $\delta_{f}$). However, the notable difference between equations (14), (15) and (9),(10) lies in accommodating the biases of section 2.2.3. This involves the inclusion of three dummy variables – $I_{Pref}$, $I_{Seq_{s}}$, and $I_{Seq_{f}}$ – interacting with the number of successes (in (14)) or failures (in (15)). $I_{Pref}$ equals 1 when the subject faces a decision with preferences over the state-space (that is, when beliefs about a dollar urn were reported) and 0 otherwise. $I_{Seq_{s}}$ ($I_{Seq_{f}}$) equals 1 when the last three balls observed in the second sequence are red (blue). By doing so, one can distinguish motivated beliefs (optimism and pessimism), sequence-related biases (gambler’s fallacy and hot-hand fallacy) and over/under inference. The presence of those biases is tested by comparing whether the estimated parameters are significantly different from their Bayesian values. Table 1 summarizes the baseline and complete models, and specifically shows the ranges of values of the different parameters, which correspond to each bias.
Baseline Model (Eqs. 12, 13) Complete Model (Eqs. 14, 15, 16)
Inference Biases
$\gamma_{s}$ or $\gamma_{f}>1$: Overinference $\alpha_{0}$ or $\beta_{0}>1$: Overinference
$\gamma_{s}$ or $\gamma_{f}\in[0,1)$: Underinference $\alpha_{0}$ or $\beta_{0}\in[0,1)$: Underinference
$\gamma_{s}$ or $\gamma_{f}<0$: Against signal $\gamma_{s}$ or $\gamma_{f}<0$: Against signal
Base-Rate Biases
$\delta_{s}$ or $\delta_{f}>1$: Base-Rate Overuse $\delta_{s}$ or $\delta_{f}>1$: Base-Rate Overuse
$\delta_{s}$ or $\delta_{f}<1$: Base-Rate Neglect $\delta_{s}$ or $\delta_{f}<1$: Base-Rate Neglect
Confirmation Biases
$\rho_{s}$ or $\rho_{f}<0$: Confirmation Bias
$\rho_{s}$ or $\rho_{f}>0$: Disconfirmation Bias
Preference-Based Biases
$\alpha_{0}+\alpha_{Pref}>1$ or $\beta_{0}+\beta_{Pref}<1$: Optimism
$\alpha_{0}+\alpha_{Pref}<1$ or $\beta_{0}+\beta_{Pref}>1$: Pessimism
Asymmetries:
$\alpha_{Pref}>\beta_{Pref}$: Good News Effect
$\alpha_{Pref}<\beta_{Pref}$: Bad News Effect
Sequence-related Biases
$\alpha_{0}+\alpha_{Seq}>1$ or $\beta_{0}+\beta_{Seq}>1$: Hot-Hand Fallacy
$\alpha_{0}+\alpha_{Seq}<0$ or $\beta_{0}+\beta_{Seq}<0$: Gambler’s Fallacy
Precision Biases
$\nu>1$: Underprecision
$\nu<1$: Overprecision
Table 1: Summary of biases for Baseline and Complete Model
4.3 The effect of each individual bias
A bias may be frequently detected across subjects and yet have only a modest quantitative effect on inference when it appears. Conversely, a less prevalent bias may generate large distortions in posterior beliefs. For this reason, beyond classifying which biases are present, I also construct a measure of the relative importance of each bias in driving deviations from the Bayesian benchmark. This measure identifies the specific effect of each bias in the relative changes of mean and variance for each subject. In order to achieve this, I compare the Bayesian expected value and Bayesian variance to bias-specific counterfactual moments that isolate the effect of one particular bias. I refer to these objects as Bias-specific expected value and Bias-specific variance. The Bayesian expected value and variance of the posterior beta distribution are given by:
$E_{n}$ $\equiv E(p|a_{n},b_{n})=\frac{a_{n}}{a_{n}+b_{n}}$ (17)
$Var_{n}$ $\equiv Var(p|a_{n},b_{n})=\frac{a_{n}b_{n}}{(a_{n}+b_{n})^{2}(a_{n}+b_{n}+1)}$ (18)
The bias-specific expected value $E_{Bias}$ and bias-specific variance $Var_{Bias}$ [^19] In the case of precision biases, $Var_{Bias}$ is already given by $\tilde{Var}_{n}$, and $E_{Bias}$ would be unaffected. are given by:
$E_{Bias}$ $=\frac{a_{bias}}{a_{bias}+b_{bias}}$ (19)
$Var_{Bias}$ $=\frac{a_{bias}b_{bias}}{(a_{bias}+b_{bias})^{2}(a_{bias}+b_{bias}+1)}$ (20)
where $a_{bias}$ and $b_{bias}$ are the hypothetical parameters of a posterior beta distribution if they were distorted by only one specific bias. [^20] Because we are looking at the effect of specific biases for each subject, it is theoretically possible that the effect of the bias in isolation is so strong that it makes $a_{bias}$ or $b_{bias}$ negative. This would make $E_{Bias}$ and $Var_{Bias}$ uninterpretable. Therefore, it is assumed that the maximum effect that a bias in isolation can have, is such that the associated parameter $a_{bias}$ or $b_{bias}$ is equal to 0 + $\varepsilon$. It is worth noting that these cases were very rare, as it would require strongly updating against the information signal. For example, suppose we want to evaluate the impact of overinference. Suppose a particular subject exhibits overinference only in successes: That is $\alpha_{0}>1$. Then, $a_{bias}=a_{Overinference}=\alpha_{0}k+a_{0}$ and $b_{bias}=b_{n}$. Equations (19) and (20) for every other bias are calculated analogously.
To capture the relative importance of each bias in inference, I compare the distance between the Bayesian and distorted expected value and variance. Namely:
$\Delta E_{Bias}$ $=|E_{n}-E_{Bias}|$ (21)
$\Delta Var_{Bias}$ $=|Var_{n}-V_{Bias}|$ (22)
Equations (21) and (22) allow us to quantify the discrepancy between the Bayesian expected value and variance, and the bias-specific measures. These discrepancies serve as indicative measures of the influence of each bias on the overall inference process.
5 Results
5.1 Population-level analysis
Table 2 reports the baseline model (equations (12) and (13)) and the complete model (equations (14), (15) and (16)) estimated at the population level. Columns (1) and (2) correspond to the baseline model, while columns (3), (4) and (5) correspond to the complete model. All regressions in Table 2 are estimated with standard errors clustered at the participant level. [^21] A negligible fraction of observations (4 out of 2640, or 0.1%) yielded slightly negative values for $a_{0}$ and $b_{0}$ due to a minor margin of error in the graphical interface. This only occurred in rare cases where the reported variance was exceptionally high and the expected value was near the extremes (0.01 or 0.99). These observations were excluded from the regressions.
Dependent variable:
$a$ posteriorposterior $b$ postpost $a$ postpost $b$ postpost VarianceVariance postpost
(1) (2) (3) (4) (5)
SuccessesSuccesses 37.234∗∗ 43.678
(11.66) (36.34)
$a$ priorprior 0.017∗∗∗ 0.016∗∗∗
(0.02) (0.02)
FailuresFailures 76.199∗∗ 67.152
(32.01) (60.956)
$b$ priorprior -0.0002∗∗∗ -0.001∗∗∗
(0.019) (0.02)
$Success:preference$ -30.889
$Success:Seq_{pos}$ 14.726
$Failures:preference$ -27.703
$Failures:Seq_{neg}$ 102.810
ConfirmationConfirmation 0.001 -0.097
(0.001) (0.09)
BayesianBayesian variancevariance 0.992
ConstantConstant 0.002∗∗
Observations 2,636 2,636 2,636 2,636 2,636
R2 0.002 0.003 0.002 0.004 0.289
Adjusted R2 0.001 0.002 0.00000 0.002 0.288
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01 relative to the Bayesian benchmark
Table 2: Baseline and complete models at the population level. Significance is with respect to Bayesian values. Clustered standard errors by participant.
As in the rest of the paper, significance is defined relative to the corresponding Bayesian benchmark rather than relative to zero (see Table 1). In the baseline specification, two biases are significant at the population level: overinference and base-rate neglect. In particular, the coefficients on successes and failures are significantly above their Bayesian benchmarks, while the coefficients on prior beliefs are significantly below one. Thus, when the data are summarized at the aggregate level, the baseline model suggests that subjects overweight the information signal and underweight prior beliefs.
Once the full set of biases is incorporated, the evidence for population-level overinference disappears, whereas base-rate neglect remains significant. This is the main comparison between the baseline and complete specifications at the aggregate level: when a richer set of biases is factored in, baseline overinference estimates appear to be partly absorbing other distortions. This interpretation is reinforced by the information criteria reported in Table 3 in Appendix C, where the complete model improves upon the baseline specification. In addition, Appendix E reports a Grether-style reduced-form specification of the same data in the traditional binary-state framework. The resulting estimates are qualitatively consistent both with the classic belief-updating literature and with the broad aggregate patterns reported here.
At the same time, Table 2 shows that the population-level equations explain little of the variation in posterior mean reports. The goodness of fit is extremely low in columns (1) to (4), for both the baseline and complete specifications. However, this should not be interpreted as the model being unable to explain subjects’ updating behavior. Rather, it reflects the fact that different subjects exhibit different distortions, which tend to average out in the population regression. Consistent with this interpretation, Table 4 in Appendix C shows that the goodness of fit of all equations improves markedly when the model is estimated separately for each participant, while Table 3 shows analogous improvements in AIC and BIC.
These results point to substantial individual heterogeneity in belief-updating distortions and motivate the individual-level analysis that follows.
5.2 Individual-level analysis
Figures 4(a) and 4(b) provide an overview of the biases detected at the individual level within the context of the two models: the baseline and the complete model. In Figure 4(a), the bar chart depicts the occurrences of statistically significant biases at the 5% level within the baseline model. The color-coded representation distinguishes biases: red means significance solely for successes, blue for failures, and green for significance in both successes and failures. Notably, overinference and base-rate overuse emerge prominently. However, Figure 4(b), illustrating individual biases in the complete model, shows a more complete picture.
Here, a broader array of biases is accounted for, with the hot-hand fallacy being the most commonly exhibited bias among the subjects. [^22] This is especially surprising given that sequence-related biases, such as the hot-hand fallacy, are tested with a reduced number of observations, which generally limits statistical power and thus makes significant findings less expected. Further comparing these models yields interesting insights. First, at the descriptive level, it becomes evident that individual heterogeneity persists across both models: All tested biases were present among subjects to a certain extent. That is, there is no bias which fully vanishes after incorporating the complete set of biases. Notably, the prevalence of the hot-hand fallacy contrasts sharply with the sparsity of confirmatory biases or underprecision among subjects. Second, moving from the baseline to the complete model substantially reduces the set of subjects who remain uncategorized.
(a) Figure 4a: Number of times a specific bias is found to be significant ($p<0.05$) in the baseline model at the individual level.
(b) Figure 4b: Number of times a specific bias is found to be significant ($p<0.05$) in the complete model at the individual level.
While the “No Bias” bar in Figure 4(a) shows that 15 individuals (17% of the sample) could not be categorized –that is, they are either too noisy to classify or statistically indistinguishable from Bayesian updaters– in the complete model (Figure 4(b)) this number falls to 2. Since the complete model tests a larger number of candidate distortions, Appendix D.2 revisits these comparisons under a model-specific multiple-hypothesis testing correction. The main conclusions remain largely unchanged: all biases continue to be present, the hot-hand fallacy is the most commonly exhibited bias, and the complete model still classifies substantially more subjects than the baseline model.
Despite these analyses showing the prevalence of different biases at the individual level, this does not tell us which biases matter most for inference. A bias may be common but quantitatively modest, or relatively less frequent but highly consequential when it appears. Thus, I apply the methodology outlined in section 4.3, to assess the relative importance of each bias.
Figures 5(a) and 5(b) examine biases within the complete model, focusing on the expected-value deviations from the Bayesian framework for each specific bias. While both figures provide insights into how important these deviations are, they take different approaches. Figure 5(a) presents the expected value deviations for each bias, irrespective of their frequency of occurrence. On the other hand, Figure 5(b) adjusts for frequency by weighting these deviations with the number of times a bias was found to be statistically significant.
Effects on Expected Value and Variance
(a) Figure 5a: Average expected-value deviations from each individual bias.
(b) Figure 5b: Expected-value deviations adjusted by bias prevalence (at $p<0.05)$.
In Figure 5(a), two biases, gambler’s fallacy and optimism, stand out with prominent expected value deviations ($\approx 0.38$ and $\approx 0.35$ distance units to the Bayesian posterior expected value respectively), which emphasizes their substantial impact on inference. In contrast, confirmation and disconfirmation bias exhibit less pronounced deviations. Remarkably, even after correcting for frequency in Figure 5(b), gambler’s fallacy and optimism retain their prominence, suggesting a clear influence on biased inference. [^23] Figures 5(b) and 6(b) should be interpreted ordinally.
Beyond these observations, Figure 5(a) and 5(b) also reveal interesting patterns. The biases exerting the most substantial influence on expected-value deviations cluster into two distinct categories: Motivated beliefs (optimism and pessimism), and biases associated with updating against the information signal, including gambler’s fallacy. Additionally, there is a notable observation regarding the hot-hand fallacy: While identified as the most common bias in Figure 4(b), its impact on expected value deviations appears relatively modest in Figure 5(a). However, when it comes to its effect on variance, the hot-hand fallacy demonstrates a notably strong influence, as can be seen in Figures 6(a) and 6(b).
Figure 6(a) and 6(b) are analogous to 5(a) and 5(b), but represent variance deviations with respect to the Bayesian framework instead. While the relative contribution of each bias appears to be quite homogeneous (Figure 6(a)), once we adjust for bias prevalence such biases (Figure 6(b)), the hot-hand fallacy shows its prominence, while optimism and pessimism come second and third respectively. This underlines the importance of motivated-belief biases in overall inference: Both in expected-value and variance deviations.
(a) Figure 6a: Average variance deviations from each individual bias.
(b) Figure 6b: Variance deviations adjusted by bias prevalence (at $p<0.05)$.
Putting it all together, the individual-level analysis shows that the complete model provides a richer account of heterogeneity than the baseline model, and identifies sequence-related and motivated-belief biases as the main sources of departures from Bayesian updating.
5.3 Co-occurrence patterns across belief biases
Biases rarely appear in isolation. To summarize heterogeneity in a compact and easily interpretable way, I study how the biases estimated in the complete model co-occur across subjects. [^24] This analysis is descriptive and exploratory: it was not pre-registered in this exact form, and it is included as a complement to the individual-level estimates. Let $j\in\{1,\dots,J\}$ index the set of bias labels considered in the complete model (e.g., over-/under-inference, base-rate distortions, motivated-belief asymmetries, sequence-related biases, (over/under)precision, and confirmation/disconfirmation). For each subject $i\in\{1,\dots,N\}$, I define a binary indicator $B_{ij}\in\{0,1\},$ which equals one if subject $i$ is classified as exhibiting bias $j$ (according to the individual-level estimation), and zero otherwise. Stacking these indicators yields an $N\times J$ subject-by-bias matrix $B$.
For each bias $j$, I compute its prevalence
$\hat{\upsilon}_{j}\;=\;\frac{1}{N}\sum_{i=1}^{N}B_{ij}$ (23)
,which is simply a subject-normalized score for Figure 4(b) (e.g. hot-hand: $\hat{\upsilon}\approx 0.44$, overinference $\hat{\upsilon}\approx 0.31$, base-rate distortions $\hat{\upsilon}\approx 0.29$ etc.). [^25] Because some biases are low-prevalence, associations involving these biases should be interpreted cautiously.
To quantify co-occurrence between two binary bias indicators $B_{\cdot j}$ and $B_{\cdot j}$, I compute the phi correlation coefficient, $\phi_{jj^{\prime}}=\mathrm{corr}(B_{\cdot j},B_{\cdot k}),$ which coincides with the Pearson correlation computed on $\{0,1\}$-valued variables. Positive values indicate that the two biases tend to be jointly exhibited, while negative values indicate that they tend to be mutually exclusive.
Figure 7: Bias co-occurrence heatmap. Each cell reports the phi correlation $\phi_{jj^{\prime}}=\mathrm{corr}(B_{\cdot j},B_{\cdot j^{\prime}})$ between two bias indicators across subjects. Biases are ordered by hierarchical clustering (average linkage) using distance $1-\phi_{jj^{\prime}}$.
I use two complementary visualizations. Figure 7 displays the full matrix of pairwise co-occurrence coefficients. The hierarchical ordering is used purely for readability, so that positively related biases appear as reddish blocks and systematic negative relationships as bluish contrasts. The dendrogram along the margins shows the resulting clustering.
Figure 8: Bias co-occurrence network. Nodes represent biases and node size is proportional to prevalence $\hat{\upsilon}_{j}$. A red edge is drawn when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=1),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=1)\}\geq 0.5$. A blue edge is drawn when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=0),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=0)\}\geq 0.5$. Edge thickness is proportional to $|\phi_{jj^{\prime}}|$.
Figure 8 provides a condensed summary of the strongest co-occurrence patterns, as a network between biases. Nodes represent biases and node size is proportional to prevalence $\hat{\upsilon}_{j}$. An edge between two biases $j$ and $j^{\prime}$ is drawn when their association is strong enough to have a transparent probabilistic interpretation. Specifically, I draw a red edge (positive co-occurrence) when observing one bias makes the other more likely than not, i.e. when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=1),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=1)\}\geq 0.5$. Conversely, I draw a blue edge (negative co-occurrence) when the reverse holds in the sense of near-mutual exclusivity, i.e. when the presence of bias $j$ makes it more unlikely that bias $j^{\prime}$ is happening ( $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=0),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=0)\}\geq 0.5$). Edge thickness is proportional to $|\phi_{jj^{\prime}}|$.
Both figures point to a small number of clear qualitative patterns. First, the two sequence-related biases exhibit a strong negative relationship: hot-hand and gambler’s fallacy tend not to be jointly exhibited. In the individual-level data, $\phi(\text{Gambler},\text{HotHand})\approx-0.36$, and the conditional probabilities reflect this asymmetry: $\Pr(\text{HotHand}=1\mid\text{Gambler}=1)\approx 0.11$ while $\Pr(\text{HotHand}=1\mid\text{Gambler}=0)\approx 0.54$. This relationship appears as a blue contrast in the heatmap and as the only prominent blue edge connecting the two sequence nodes in the network. [^26] Note that even though this strong negative correlation is highly intuitive, these biases are (in theory) not necessarily mutually exclusive by definition, as there could be a success-failure asymmetry.
Second, and perhaps more importantly, this divide permeates through the rest of the bias set. The resulting graph from Figure 8 –which highlights the strongest co-occurrence relationships– has a roughly path-like form, with the sequence-related biases (hot-hand and gambler’s fallacy) acting as a hub that separates two main network components. On one side, prior-based distortions and updating against the signal are closely connected, with the gambler/against link and its connections to base-rate overuse forming a compact cluster. On the other side, a set of “jumping-to-conclusions” distortions appear jointly: overprecision and optimism are directly connected with the hot-hand fallacy, while overinference, pessimism and motivated belief asymmetries are only strongly connected to the hot-hand effect through overprecision. The same component also connects to confirmation bias. [^27] The presence of confirmation, and underprecision in this component should be interpreted with caution given their relatively low prevalence, but the position of confirmation bias in the network is consistent with appearing alongside strong updating and concentrated posterior beliefs.
Finally, several pairs of biases that correspond to opposing distortions are negatively related in the co-occurrence matrix. Beyond the hot-hand/gambler opposition discussed above, base-rate neglect and base-rate overuse are negatively related, $\phi(\text{BRN},\text{BRO})\approx-0.24$, and overinference is negatively related to underinference, $\phi(\text{Overinfer},\text{Underinfer})\approx-0.21$ (etc.) At the same time, some conceptually related biases co-occur strongly: e.g. gambler’s fallacy is positively associated with updating against the signal ($\phi\approx 0.36$), and the hot hand with overprecision ($\phi\approx 0.17$).
Overall, the co-occurrence structure suggests that biases cluster into a small number of recurring patterns rather than appearing independently, and further highlights the importance of sequence-related biases. Because this analysis is partly exploratory and some biases are low-prevalence, I treat these patterns as descriptive regularities rather than as definitive typologies.
6 Concluding remarks
This paper develops a unified framework for separately identifying multiple belief-updating biases that may otherwise be empirically confounded. The key methodological idea is to study belief updating using prior and posterior belief distributions over a continuous state space, rather than point beliefs alone. This allows a broader range of conflicting distortions to be brought into the same structural framework and tested in a laboratory setting using a novel belief-elicitation method.
The results show that accounting for a richer set of biases changes the interpretation of updating behavior in important ways. At the population level, overinference and base-rate neglect are both significant in the baseline specification, but once the complete model is estimated, only base-rate neglect remains significant. At the individual level, the analysis uncovers substantial heterogeneity: under the complete specification, all tested bias types are present to some extent in the data, and more subjects are categorized as exhibiting some type of bias. Motivated-belief biases and sequence-related biases emerge as the main drivers of biased inference, while confirmatory distortions are comparatively scarce. Moreover, biases do not arise independently. Rather, they exhibit systematic co-occurrence patterns that suggest a small number of recurring configurations, with sequence-related distortions appearing central in the overall structure.
Taken together, these findings suggest that studying belief-updating biases one at a time may give an incomplete picture of the underlying mechanisms shaping inference. More generally, a framework that separates multiple competing distortions can be useful for researchers interested in the interaction between biases, as well as for work seeking belief-based explanations of broader behavioral phenomena, such as political polarization, investment behavior, or the disposition effect.
Several directions for future research remain open. While this paper takes detailed care in separating behavior stemming from different biases, it treats such biases as systematic departures from Bayesian updating and does not attempt to identify the deeper cognitive mechanisms through which they arise. In this direction, Bordalo et al. (2025) study how some biases become more prominent when salient features of information become more relevant. Likewise, Bordalo et al. (2023) emphasize the role of memory and recall in generating well-known cognitive distortions, while Graeber et al. (2024) show how different forms of information –such as narratives versus statistical evidence– affect belief updating. Exploring these mechanisms in settings where biases can already be separately identified may help provide a more complete account of which updating distortions matter most, and of the underlying primitives that generate biased behavior.
Appendix
Appendix A Theoretical derivations
This section shows the derivation of equations of posterior parameters in section 2. I present two cases: Proposition 1 characterizes the Bayesian benchmark under standard beta-binomial conjugacy (eq.(3)), and Proposition 2 introduces the biases of the baseline model (eq.(4) and (5)). The rest of the derivations for the remaining biases are analogous to these cases. Once the distorted likelihood and prior are given, the proof follows the exact same steps.
Proposition 1.
Given likelihood equation (1) and prior equation (2) of section 2.1., a Bayesian agent updates her beliefs such that her posterior distribution of $p$: $\pi(p|a_{n},a_{n})$ is beta distributed with parameters $a_{n},b_{n}$ such that:
$a_{n}=k+a_{0}\qquad b_{n}=n-k+b_{0}$
Proof.
Let us apply Bayes’ Theorem given likelihood eq.(1) and prior eq.(2). This yields:
$\pi(p|s_{1}...s_{n},a_{0},b_{0})$ $=\frac{L(p|s_{1}...s_{n})\pi(p|a_{0},b_{0})}{\int_{p=0}^{1}\big(L(p|s_{1}...s_{n})\pi(p|a_{0},b_{0})\big)dp}$
$=\frac{\binom{n}{k}p^{k+a_{0}-1}(1-p)^{n-k+b_{0}-1}/B(a_{0},b_{0})}{\int_{p=0}^{1}\Big(\binom{n}{k}p^{k+a_{0}-1}(1-p)^{n-k+b_{0}-1}/B(a_{0},b_{0})\Big)dp}$
$=\frac{p^{k+a_{0}-1}(1-p)^{n-k+b_{0}-1}}{B(a_{0}+k,b_{0}+n-k)}$
Which is itself the probability density function of a beta distribution with parameters ($a_{0}+k,b_{0}+n-k$). Therefore, the posterior Bayesian distribution is beta distributed with parameters ($a_{n},b_{n}$) as defined in Proposition 1.
Proposition 2.
Given likelihood $\tilde{L}(p|s_{1}...s_{n})=(L(p|s_{1}...s_{n}))^{\gamma}$ and prior $\tilde{\pi}(p)=(\pi(p))^{\delta}$ of section 2.1., where $\gamma,\delta$ indicate deviations from Bayesian updating, a non-Bayesian agent follows a posterior beta distribution with parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
$\tilde{a}_{n}=\gamma k+\delta(a_{0}-1)+1$
$\tilde{b}_{n}=\gamma(n-k)+\delta(b_{0}-1)+1$
Proof.
Applying Bayes theorem yields:
$\tilde{\pi}(p|s_{1}...s_{n},a_{0},b_{0})$ $=\frac{(L(p|s_{1}...s_{n}))^{\gamma}(\pi(p))^{\delta}}{\int_{p=0}^{1}\big((L(p|s_{1}...s_{n}))^{\gamma}(\pi(p))^{\delta}\big)dp}$
$=\frac{\big[\binom{n}{k}\big]^{\gamma}p^{\gamma k}(1-p)^{\gamma(n-k)}\big(1/B(a_{0},b_{0})\big)^{\delta}p^{\delta(a_{0}-1}(1-p)^{\delta(b_{0}-1)}}{\int_{p=0}^{1}\Big(\big[\binom{n}{k}\big]^{\gamma}p^{\gamma k}(1-p)^{\gamma(n-k)}\big(1/B(a_{0},b_{0})\big)^{\delta}p^{\delta(a_{0}-1}(1-p)^{\delta(b_{0}-1)}\Big)dp}$
$=\frac{p^{\gamma k+\delta(a_{0}-1)}(1-p)^{\gamma(n-k)+\delta(b_{0}-1)}}{\int_{p=0}^{1}\Big(p^{\gamma k+\delta(a_{0}-1)}(1-p)^{\gamma(n-k)+\delta(b_{0}-1)}\Big)dp}$
$=\frac{p^{\gamma k+\delta(a_{0}-1)}(1-p)^{\gamma(n-k)+\delta(b_{0}-1)}}{B\Big(\gamma k+\delta(a_{0}-1)+1,\gamma(n-k)+\delta(b_{0}-1)+1\Big)}$
This is the probability density function of a beta distribution with parameters $\Big(\gamma k+\delta(a_{0}-1)+1,\gamma(n-k)+\delta(b_{0}-1)+1\Big)$. Therefore, the posterior distribution of an agent that exhibits inference bias $\gamma$, and base-rate bias ($\delta$) is beta distributed with parameters ($\tilde{a}_{n},\tilde{b}_{n}$) as defined in Proposition 2. ∎
Appendix B Incentivizing moments of beta distributions
The scoring rules follow Schlag and Van der Weele (2013) with a slight modification. In particular, random realizations of the Bayesian posterior distributions are taken as the random draws.
Incentivizing the mean.
Let $\tilde{m}$ be the reported mean of the agent’s posterior beta distribution $\pi(p|\tilde{a}_{n},\tilde{b}_{n})$, and let $d$ be a random draw of the Bayesian posterior beta distribution $\pi(p|a_{n},b_{n})$. Then the Quadratic Scoring Rule is given by $g_{QSR}(\tilde{m},d)=-(\tilde{m}-d)^{2}$. Let $A,B$ be the boundaries of the state-space $\Omega$, and $M$ be any arbitrary amount of money. [^28] In the experiment $A=1$, $B=99$ and $M=25/3$ cents for each report (that is a total maximum of 10 euros) Then, the randomized quadratic scoring rule is given by the following lottery:
$\tilde{g}_{QSR}(\tilde{m},d)=l\Bigg(M,0;1+\frac{g_{QSR}(\tilde{m},d)}{(B-A)^{2}}\Bigg)$
Incentivizing the variance.
Let $\tilde{v}$ be the the reported variance of the agent’s posterior beta distribution. In order to elicit the variance consider two random draws of the agent’s posterior beta distribution $\pi(p|\tilde{a}_{n},\tilde{b}_{n})$. Namely, $d_{1}$ and $d_{2}$. Then the variance scoring rule is given by $g_{v}(\tilde{v},d_{1},d_{2})=-\Big(\tilde{v}-\frac{1}{2}(d_{1}-d_{2})^{2}\Big)^{2}$. Applying randomisation, the randomized variance scoring rule yields:
$\tilde{g}_{v}(\tilde{v},d_{1},d_{2})=l\Bigg(M,0;\frac{g_{v}+\frac{1}{4}(B-A)^{4}}{\frac{1}{4}(B-A)^{4}}\Bigg)$
Appendix C Extra tables and figures
Model AIC(pop.) BIC(pop.) AIC(ind.) BIC(ind.)
Baseline model
      Eq. (12) 49858.50 49876.13 269.42 273.61
      Eq. (13) 51038.89 51056.52 265.07 269.27
Complete model
      Eq. (14) 49863.44 49898.70 267.81 276.20
      Eq. (15) 51041.91 51077.17 254.25 262.65
      Eq. (16) -14672.77 -14655.14 -187.01 -182.82
Table 3: Information Criteria Comparison at population and individual level
Model $R^{2}$ Population Mean $R^{2}$ Ind. Mean adj. $R^{2}$ Ind.
Baseline model
      Eq. (12) 0.002 0.565 0.534
      Eq. (13) 0.003 0.535 0.501
Complete model
      Eq. (14) 0.002 0.658 0.589
      Eq. (15) 0.004 0.720 0.664
      Eq. (16) 0.292 0.321 0.297
Table 4: $R^{2}$ Comparison for Baseline and Complete model at population vs. individual level
Appendix D Robustness analyses
D.1 Robustness to Measurement Error in Uncertainty Reports
Uncertainty is elicited through a continuous slider and may therefore be measured with classical measurement error. To assess whether such measurement error affects the bias classification, I borrow the noise-injection logic from the SIMEX (Simulation Extrapolation) method Cook and Stefanski (1994) for parametric models, and apply it to both the prior and posterior reported variance.
Let $\mu$ denote the reported belief mean (be it prior or posterior) and $V$ the reported variance. Because beliefs are defined on the bounded support $[0,1]$, reported variances must satisfy
$0<V<V_{\max}(\mu),\qquad V_{\max}(\mu)=\mu(1-\mu)$
Note that this restriction follows from the support of the distribution and does not impose unimodality of the implied beta distribution; bimodal beta distributions remain feasible whenever the reported variance is sufficiently close to the upper bound. To model reporting noise while preserving feasibility, I work with the normalized variance ratio
$u=\frac{V}{V_{\max}(\mu)}\in(0,1)$
and apply the logit transformation $z=\log\!\left(\frac{u}{1-u}\right)\in\mathbb{R}$.
Measurement error is assumed to be classical and additive on this scale:
$z^{\text{rep}}=z^{\star}+\sigma_{z}\varepsilon\qquad\varepsilon\sim\mathcal{N}(0,1)$
where $z^{\star}$ denotes the latent (noise-free) report and $\sigma_{z}$ is the standard deviation of reporting noise on the logit-normalized variance scale.
To implement this robustness exercise, I simulate additional noise on the logit-normalized variance scale. For each $\lambda\in\Lambda=\{0.1,0.25,0.5\}$ and each observation, the perturbed report is given by
$z^{(\lambda)}=z^{\text{rep}}+\sqrt{\lambda}\,\sigma_{z}\,\varepsilon^{(\lambda)}\qquad\varepsilon^{(\lambda)}\sim\mathcal{N}(0,1)$ (24)
where $\lambda$ indexes the severity of the perturbation. Thus, $\lambda=0.1$, $\lambda=0.25$, and $\lambda=0.5$ correspond to adding noise with variance equal to $10\%$, $25\%$, and $50\%$ of the estimated baseline measurement-error variance, respectively. This range is intended to capture progressively stronger perturbations of the reported variance while remaining within a plausible range for slider misreporting.
A baseline value $\sigma_{z}=\hat{\sigma}_{z}$ is calibrated from the data as the residual standard deviation from a regression of reported logit-normalized variances on their Bayesian benchmark variances on the same scale. Mapping back via $u^{(\lambda)}=\text{logit}^{-1}(z^{(\lambda)})$ yields simulated feasible variances (both for prior and posterior measurements)
$V^{(\lambda)}=u^{(\lambda)}V_{\max}(\mu)$
For each $\lambda$, the simulation is repeated $200$ times and the adjusted variances $V^{(\lambda)}$ are averaged across repetitions. The procedure is applied to both prior and posterior variance reports. For each simulated variance, the implied beta-distribution parameters are recomputed and all model equations are re-estimated using these noise-adjusted values.
This section replicates Figures 4(a), 4(b), 5(b) and 6(b) in the main text for the largest simulated noise level, $\lambda=0.5$.
(a) Figure 9a: Number of times a specific bias is found to be significant ($p<0.05$) in the baseline model at the individual level after noise injection of $\lambda=0.5$ in reported variances.
(b) Figure 9b: Number of times a specific bias is found to be significant ($p<0.05$) in the complete model at the individual level after noise injection of $\lambda=0.5$ in reported variances.
As Figures 9(a) and 9(b) show, many of the main qualitative conclusions remain unchanged even under this conservative level of noise injection. Every bias continues to be present in the complete model, the complete model still categorizes a substantially larger share of subjects than the baseline model, and hot-hand effect remains the most commonly exhibited bias. Moreover, Figures 10(a) and 10(b) show that sequence-related biases (gambler’s fallacy and hot-hand behavior) and motivated beliefs (especially optimism) continue to account for most of the inference in expected-value and variance deviations. The main change is that overinference becomes less prevalent, while underinference becomes substantially more prominent.
(a) Figure 10a: Expected-value deviations with respect to the Bayesian framework (adjusted by bias prevalence) after noise injection of $\lambda=0.5$.
(b) Figure 10b: Variance deviations with respect to the Bayesian framework (adjusted by bias prevalence) after noise injection of $\lambda=0.5$.
D.2 Individual bias frequency corrected for multiple hypothesis testing
As the individual-level analysis in Section 5.2 classifies subjects on the basis of multiple coefficient tests, a natural concern is that some detected biases may reflect false positives. This concern is particularly relevant when comparing the baseline and complete models, since the baseline model tests four hypotheses per subject, whereas the complete model tests twelve. To address this issue, this appendix repeats the individual-level analysis after applying a Bonferroni correction within each model. Accordingly, the significance threshold is set to $p<0.05/4=0.0125$ for the baseline model, and to $p<0.05/12\approx 0.0042$ for the complete model. This section replicates Figures 4(a), 4(b), 5(b) and 6(b) under these corrected thresholds.
(a) Figure 11a: Number of times a specific bias is found to be significant ($p<0.0125$) in the baseline model at the individual level after multiple hypothesis testing correction.
(b) Figure 11b: Number of times a specific bias is found to be significant ($p<0.0042$) in the complete model at the individual level after multiple hypothesis testing correction.
(a) Figure 12a: Expected-value deviations with respect to the Bayesian framework (adjusted by bias prevalence at $p<0.0042$).
(b) Figure 12b: Variance deviations with respect to the Bayesian framework (adjusted by bias prevalence at $p<0.0042$)
The main qualitative conclusions remain largely unchanged: Every bias continues to be present in the complete model, and the hot-hand fallacy continues to be the most commonly exhibited bias (See Figures 11(a) and 11(b)). Also, sequence-related biases remain the main drivers of distorted inference, with gambler’s fallacy continuing to play a leading role for expected-value deviations and hot-hand behavior for variance deviations (See Figures 12(a) and 12(b)).
At the same time, the correction naturally reduces the number of detected significant biases and increases the number of unclassified subjects in both models. Nevertheless, the complete model still classifies substantially more subjects than the baseline model. In particular, the number of subjects in the “No Bias” category falls from $31$ under the baseline model to $20$ under the complete model (a $\approx 35\%$ reduction). Finally, it is important to note that, motivated-belief biases –especially optimism and pessimism– become relatively less prevalent under the corrected thresholds.
Appendix E A Grether-style reduced-form model
This section connects the main continuous state-space analysis to the classic reduced-form literature on belief updating in the spirit of Grether (1980). To do so, I approximate the continuous state space of the model by a binary state space and estimate a Grether-style regression using the experimental data. This exercise is intended to show that applying the usual models to this experimental data recovers patterns that can be related both to the classic literature and to the baseline findings of the main model.
Let the state space $\Theta=\{L,H\}$ be a binary partition of the original state-space $\Omega=(0,1)$, where state $H$ corresponds to higher success probabilities and state $L$ to lower success probabilities.
Let $q_{0}=\Pr(H)$ denote the prior belief assigned to the high state, and let $q_{n}=\Pr(H\mid k,n)$ denote the corresponding posterior belief after observing $k$ successes in $n$ Bernoulli trials. Conditional on state $\theta\in\{L,H\}$, the probability of observing $k$ successes is
$\Pr(k\mid\theta)=\int_{p\in\theta}\binom{n}{k}p^{k}(1-p)^{n-k}f(p\mid\theta)\,dp$
where $f(p\mid\theta)$ denotes the prior density over $p$ conditional on state $\theta$.
The associated log-likelihood ratio is therefore
$\ell(k,n)=\log\left(\frac{\Pr(k\mid H)}{\Pr(k\mid L)}\right)$
Under Bayesian updating, posterior log-odds satisfy
$\log\left(\frac{q_{n}}{1-q_{n}}\right)=\log\left(\frac{q_{0}}{1-q_{0}}\right)+\ell(k,n)$
A Grether-style reduced form allows the prior and the likelihood ratio to receive separate weights in the updating equation. I therefore estimate
$\log\left(\frac{\tilde{q}_{n}}{1-\tilde{q}_{n}}\right)=\beta_{0}+\beta_{P}\log\left(\frac{q_{0}}{1-q_{0}}\right)+\beta_{L}\,\ell(k,n)+\varepsilon$ (25)
where $\tilde{q}_{n}$ denotes the reported posterior probability of the high state. The Bayesian benchmark is given by
$\beta_{0}=0\qquad\beta_{P}=1\qquad\beta_{L}=1$
Values $\beta_{P}<1$ indicate base-rate neglect, while $\beta_{P}>1$ indicate base-rate overuse. Similarly, $\beta_{L}<1$ indicates underinference, $\beta_{L}>1$ indicates overinference, and $\beta_{L}<0$ corresponds to inference against the signal, as in classic Grether (1980).
To implement this reduced-form exercise, I binarize the continuous state space at $p=0.5$, so that the high state is $H=\{p\geq 0.5\}$ and the low state is $L=\{p<0.5\}$. For each elicited prior and posterior beta distribution, I compute the implied probability assigned to the high state, $q=\Pr(p\geq 0.5),$ and use these probabilities as the belief objects entering the Grether-style regression. The evidence term is computed as the Bayesian log-odds update implied by the observed signal under the same binary partition, $\log\!\left(\frac{q_{n}^{B}}{1-q_{n}^{B}}\right)-\log\!\left(\frac{q_{0}}{1-q_{0}}\right),$ which is equivalent to $\ell(k,n)$ under this partition, where $q_{0}$ is the prior probability of the high state and $q_{n}^{B}$ is the Bayesian posterior probability of that state.
Dependent variable:
Log-odds reported posterior
Prior log-odds 0.681∗∗∗
Bayesian log-likelihood ratio 2.196∗∗∗
Constant 0.303∗∗∗
Observations 2,635
R2 0.559
Adjusted R2 0.559
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01
Table 5: Grether-style model at the population level. Significance is with respect to Bayesian values. Clustered standard errors by participant.
The Grether-style reduced-form estimates in Table 5 are first and foremost consistent with the baseline analysis of the paper. At the population level, the prior-weight coefficient lies below the Bayesian benchmark of one ($\hat{\beta}_{P}=0.681$), indicating base-rate neglect, while the evidence-weight coefficient lies above one ($\hat{\beta}_{L}=2.196$), indicating overinference on average. Thus, once the data are forced into a traditional binary-state reduced form, the main qualitative pattern resembles that of the baseline model: subjects appear to underweight prior information and overweight new evidence.
At the same time, these estimates remain meaningfully connected to the classic reduced-form literature on belief updating. Base-rate neglect is the standard qualitative finding in the Grether tradition, and the presence of overinference, while less canonical than underinference (Benjamin, 2019), is by no means implausible. More recent contributions also document environments in which subjects overweight new signals rather than underreact to them (Augenblick et al., 2025b; Gonçalves et al., 2025). In this sense, the Grether-style exercise places the data in a language that is comparable to the traditional literature, while still recovering the same broad reduced-form distortions highlighted by the baseline model.
The individual-level estimates (see Figure 13) likewise show substantial heterogeneity, again in a way that is closely aligned with the baseline specification. Overinference is the most prevalent reduced-form pattern, followed by base-rate neglect, while underinference and updating against the signal are also present for a non-negligible subset of participants and base-rate overuse is essentially absent.
Figure 13: Number of times a bias is found to be significant ($p<0.05$) in the Grether-style model at the individual level
Appendix F Experiment instructions
In this study, you will be asked to complete 30 guessing tasks. For each guessing task you have to make 2 related guesses. At the beginning of each guessing task, there is always a pool of 99 URNS, each containing 100 BALLS. Some balls in the urns are red, and some are blue. Each one of these urns contains a different percentage of red balls. For example, in Urn 1 there is only one red ball and 99 blue (1% of the balls are red), in Urn 2 there are only two red balls and 98 blue (2% of the balls are red). This continues until Urn 99 where 99 balls are red and one is blue. (See picture below).
For each guessing task, out of these 99 urns, one of them (say Urn X) has been selected at random. Each urn has the same chances of being selected from the pool. That is, you do not know how many of the balls are red and how many of them are blue in the selected urn. All combinations are possible. (See picture below).
Your task is to guess what the percentage of red balls in the selected urn is. In each one of the 30 guessing tasks, this process will be repeated with a new urn.
Bear in mind: Whenever a new urn is selected, it is always drawn from the same pool of 99 urns (WITH REPLACEMENT). This means that the same urn can be chosen either once or multiple times. The urns have letters (or combinations of letters) on top of them. This is just to highlight the fact that when a new letter(or combination of letters) is on top of the urn, it means that a new urn has been drawn.
The guessing task
This section will explain how each one of the 30 guessing tasks works. Initially, as explained above, the computer has randomly selected, with equal probability, ONE out of the 99 urns. Remember that initially, you know nothing about the content of this urn. Once a given urn is selected, you will be given some information about the urn to help you make your guess. First, you will see a sequence of balls which have been randomly drawn from the urn. Each ball in the selected urn has the same chances of being drawn, and each of these draws is done WITH REPLACEMENT. This means that after a ball has been drawn and taken out of the urn, it is immediately replaced with one of the same color. The urn will always have the same 100 balls. In order to help you make your guesses, two sequences of draws will be made from each urn.
  • 1st sequence of draws: Either one, two or three balls from the urn are selected at first. After this selection, you will have to make your first guess. For your first guess you will answer two questions:
    1. 1.
      What percentage of red balls do you expect the selected urn to have?
    2. 2.
      What is your uncertainty level about this percentage?
    Example: Guessing Task A
  • 2nd sequence of draws: The second draw follows very similar rules to the first one. In this case three, five or seven balls from THE SAME URN are selected with replacement. After this selection, you will have to make your second guess. Once again, for your second guess you will answer the same two questions:
    1. 1.
      What percentage of red balls do you expect the selected urn to have?
    2. 2.
      What is your uncertainty level about this percentage?
In order to further help you with your guesses a dynamic graph of your choice will be provided. Please watch the following video (next screen) to understand how this works.
——————————————[Page Break]—————————————
——————————————[Page Break]—————————————
Your Payment
You can earn up to €29,70 in this experiment. In particular, your payment is broken down as follows:
  • You will receive €5 for taking the time to complete this experiment.
  • You will receive up to €10 for your responses related to the guessing tasks.
    How much of this amount (€10) you receive depends on the actual percentage of red balls in the selected urn. You can get money for EVERY SINGLE ONE of your guesses.
    The payment rule we use, is optimized so that in order to maximize your expected payoff, you should ALWAYS give your best estimate of the percentage of red balls in the selected urn. In the same manner, the payment rule we use, is also optimized so that in order to maximize your expected payoff, you should ALWAYS give your best estimate of your uncertainty level.
  • You can receive up to €14,70 as extra payment.
    This extra payment is divided across the 30 guessing tasks. Whether or not a guessing task has an extra payment attached depends on whether a dollar urn has come up. (See picture below. Urn A is in this example, a dollar urn). In particular, a dollar urn will come up randomly in 15 of the 30 guessing tasks.
    If a dollar urn comes up, you will receive as many cents as red balls the selected dollar urn has. For example, if the selected dollar urn has 50 red balls you will get 50 cents.
At the end of the experiment, you will be informed about the number of red balls in each urn and your total payment. If you want to know more about the details of the payment rule, you can let me know after the experiment or write an email to p.gonzalezfernandez@maastrichtuniversity.nl
——————————————[Page Break]—————————————
Get Familiar with the Tools
Before you answer the comprehension questions you have the chance to get familiar with the guesses, the payment and the sliders with a trial guessing task.
——————————————[Page Break]—————————————